# Let LLMs Take on the Latest Challenges ! A Chinese Dynamic Question Answering Benchmark

### 29 Feb 2024 v1

@@ Abstract
How to better evaluate the capabilities of Large Language Models (LLMs) is the focal point and hot topic in current LLMs research. Previous work has noted that due to the extremely high cost of iterative updates of LLMs, they are often unable to answer the latest dynamic questions well. To promote the improvement of Chinese LLMs’ ability to answer dynamic questions, in this paper, we introduce CDQA, a Chinese Dynamic QA benchmark containing question-answer pairs related to the latest news on the Chinese Internet. We obtain high-quality data through a pipeline that combines humans and models, and carefully classify the samples according to the frequency of answer changes to facilitate a more fine-grained observation of LLMs’ capabilities. We have also evaluated and analyzed mainstream and advanced Chinese LLMs on CDQA. Extensive experiments and valuable insights suggest that our proposed CDQA is challenging and worthy of more further study. We believe that the benchmark we provide will become the key data resource for improving LLMs’ Chinese question-answering ability in the future.
@@ 3.1 Experiment Setup
We experiment with a series of baseline models pretrained with Chinese data, including representative OpenAI’s ChatGPT (gpt-3.5-turbo-1106) OpenAI (2022) and GPT-4 (gpt-4-1106-preview) OpenAI (2023), open-sourced Chinese-oriented models such as Internlm-20B-Chat InternLMTeam (2023), Aquila2-34B-Chat BAAI (2023), Yi-34B-Chat 01-ai (2023), Deepseek-67B-Chat DeepSeek-AI (2024), Qwen-72B-Chat Bai et al. (2023). In the close-book scenario, we only use the standalone LLM in answering questions with zero-shot and few-shot settings. For the open-book scenario, search engines are used for retrieving question-related results on the Internet and then fed into LLMs for reading.
@@ 3.2 Results and Analyses
Nevertheless, model sizes and training data are both fundamental for these prompts to work. In Figure 5, not every model improves with CoT or RaR compared to Vanilla prompt. For example, Deepseek-34B-Chat’s and Qwen-72B-Chat’s performance decrease in CoT and RaR; ChatGPT and Internlm-20B-Chat prefer RaR while GPT-4 and Yi-34B-Chat prefer CoT for larger gains; Aquila2-34B-Chat is robust to all prompt types.
@@
For a more detailed comparison between search engines across all baselines, we use vanilla prompt since CoT performs better in LLMs and zero-shot setting with Top-10 searched results from different search engines. In Figure 7, searched results from Google consistently outperforms Bing among all baseline models, which indicates that the Google currently provides the most useful retrieved evidence for question answering about Chinese news.
@@ 5 Conclusion
The creation of CDQA addresses the urgent need for the evaluation of Chinese LLMs, thereby improving LLM-driven applications for Chinese users. Given the cultural influences in LLMs’ training data, it is our aspiration that CDQA will foster development in various capabilities of LLMs, particularly within Chinese contexts. While CDQA progresses further with a semi-automatic generation pipeline with more data than FreshQA, we acknowledge that it is far from a perfect LLM evaluation. Other critical dimensions, including tool learning, LLMs safety, and robustness, remain to be explored. However, we believe that our constructed CDQA and the series of insights obtained based on it will provide valuable resources and guidance for subsequent research on Chinese LLMs.

### 02 Mar 2024 v2

@@ Abstract
How to better evaluate the capabilities of Large Language Models (LLMs) is the focal point and hot topic in current LLMs research. Previous work has noted that due to the extremely high cost of iterative updates of LLMs, they are often unable to answer the latest dynamic questions well. To promote the improvement of Chinese LLMs’ ability to answer dynamic questions, in this paper, we introduce CDQA, a Chinese Dynamic QA benchmark containing question-answer pairs related to the latest news on the Chinese Internet. We obtain high-quality data through a pipeline that combines humans and models, and carefully classify the samples according to the frequency of answer changes to facilitate a more fine-grained observation of LLMs’ capabilities. We have also evaluated and analyzed mainstream and advanced Chinese LLMs on CDQA. Extensive experiments and valuable insights suggest that our proposed CDQA is challenging and worthy of more further study. We believe that the benchmark we provide will become one of the key data resources for improving LLMs’ Chinese question-answering ability in the future.
@@ 3.1 Experiment Setup
We experiment with a series of baseline models pretrained with Chinese data, including representative OpenAI’s ChatGPT (gpt-3.5-turbo-1106) OpenAI (2022) and GPT-4 (gpt-4-1106-preview) OpenAI (2023), open-sourced Chinese-oriented models such as Internlm-20B-Chat InternLMTeam (2023), Aquila2-34B-Chat BAAI (2023), Yi-34B-Chat 01-ai (2023), Deepseek-67B-Chat DeepSeek-AI (2024). In the close-book scenario, we only use the standalone LLM in answering questions with zero-shot and few-shot settings. For the open-book scenario, search engines are used for retrieving question-related results on the Internet and then fed into LLMs for reading.
@@ 3.2 Results and Analyses
Nevertheless, model sizes and training data are both fundamental for these prompts to work. In Figure 5, not every model improves with CoT or RaR compared to Vanilla prompt. For example, Deepseek-34B-Chat’s performance decrease in CoT and RaR; ChatGPT and Internlm-20B-Chat prefer RaR while GPT-4 and Yi-34B-Chat prefer CoT for larger gains; Aquila2-34B-Chat is robust to all prompt types.
@@
For a more detailed comparison between search engines across all baselines, we use vanilla prompt since CoT performs better in LLMs and zero-shot setting with Top-10 searched results from different search engines. In Figure 7, searched results from Google slightly outperforms Bing among all baseline models, which indicates that the Google currently provides the most useful retrieved evidence for question answering about Chinese news.
@@ 5 Conclusion
The creation of CDQA addresses the urgent need for the evaluation of Chinese LLMs, thereby improving LLM-driven applications for Chinese users. Given the cultural influences in LLMs’ training data, it is our aspiration that CDQA will foster development in various capabilities of LLMs, particularly within Chinese contexts. While CDQA progresses further with a semi-automatic generation pipeline with more data than FreshQA, we acknowledge that it is far from a perfect LLM evaluation. Other critical dimensions, including tool learning, LLMs safety, and robustness, remain to be explored. However, we believe that our constructed CDQA and the series of insights obtained based on it will provide valuable resources and guidance for subsequent research on Chinese LLMs. In the future, we will conduct more in-depth analyses of the capabilities of LLMs based on CDQA and investigate how to enhance the LLMs’ ability to handle dynamic questions. This will empower LLMs to better cope with the complex and ever-changing real-world application environments.